OpenStudy (nuccioreggie):

negative eighty minus positive ninety-nine equals what value negative one hundred seventy nine negative nineteen positive nineteen positive one hundred seventy nine

OpenStudy (nuccioreggie):

@kenzieloo

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf -80-(+99)=? }\) \(\LARGE\color{blue}{ \bf -80-(+99)= }\) \(\LARGE\color{blue}{ \bf -80-99= }\) \(\LARGE\color{blue}{ \bf -179 }\)

OpenStudy (nuccioreggie):

@akshay1234

OpenStudy (anonymous):

-179

OpenStudy (nuccioreggie):

how did u get that